OpenStudy (anonymous):

prove sec^2θ/over/sec^2θ-1=csc^2

OpenStudy (anonymous):

it helps to start with \[\sec^2(x)-1=\tan^2(x)\] giving you \[\frac{\sec^2(x)}{\tan^2(x)}\]

OpenStudy (anonymous):

then you have \[\frac{\frac{1}{\cos^2(x)}}{\frac{\sin^2(x)}{\cos^2(x)}}=\frac{1}{\sin^2(x)}=\csc^2(x)\]